# TinyML NLP Approach for Semantic Wireless Sentiment Classification

### v1

@@ Abstract
Natural Language Processing (NLP) operations, such as semantic sentiment analysis and text synthesis, may often impair users privacy and demand significant on device computational resources. Centralized learning (CL) on the edge offer alternative energy efficient approach, yet requires the collection of raw information, which affects the users privacy. While Federated learning (FL) preserves privacy, it requires high computational energy on board tiny user devices. We introduce split learning (SL) as an energy efficient alternative, privacy preserving tiny machine learning (TinyML) scheme and compare it to FL and CL in the presence Rayleigh fading and additive noise. Our results show that SL reduces processing power and CO2 emissions while maintaining high accuracy, whereas FL offers a balanced compromise between efficiency and privacy. Hence, this study provides insights into deploying energy-efficient, privacy-preserving NLP models on edge devices.
@@ I Introduction
Artificial Intelligence (AI) has gained widespread adoption, with applications ranging from text and image classification to text and image generation. A key area of growth is NLP, which powers virtual assistants and human language classification systems. Large language models (LLMs), such as OpenAI’s GPT series [1], Gemini [2], and BERT, built on Transformer architectures [3], have transformed NLP by enabling machines to perform complex language tasks with high accuracy. However, these models require significant computational resources for both training and inference, posing challenges for deployment on resource-constrained devices.
@@
LLMs demand substantial storage and processing power. The high dimensionality of language data and large vocabularies further increase the computational burden [4]. Data privacy is another significant concern, as training robust models often requires diverse datasets, which can raise issues when handling sensitive information, such as in healthcare or government sectors. Additionally, deploying models on edge devices introduces communication challenges. Wireless data transmission over channels like WiFi is susceptible to noise, fading, limited bandwidth and unstable connections, which ,ay affect communication efficiency.
@@
TinyML mixed models have evolved to address the above concerns. For instance, model compression techniques, such as pruning [5], quantization [6], and knowledge distillation [7], help reduce model sizes and computational requirements. TinyBERT [8] and MobileBERT [9], for example, utilize knowledge distillation to create smaller models, while quantization techniques, like in LLaMA [10], enable efficient execution on low-resource devices. FL offers a decentralized model training approach, allowing multiple users to collaboratively train a global model while keeping their data local. Although, this method addresses data privacy concerns [11], it might be power hungry for tiny device batteries. To this end, SL [12] divides the model training process between users and a server. Users transmit activations from the initial model layers, reducing computational load and enhancing data privacy.
In this paper, we propose two frameworks to address computational challenges and privacy concerns in text emotion classification: one based on FL and the other on SL. First, we focus on semantic extraction by utilizing quantization to significantly reduce model size, enabling deployment on resource-limited devices. Second, we emphasize privacy preservation by ensuring that data remains on the device while transmitting only the necessary information for model training. The proposed methods prove roubustness in challenging channel conditions while reducing the overall energy consumption and CO2 emissions, contributing to TinyML initiatives.
@@
The remainder of this paper is structured as follows: Section II covers the system design and experimental setup, including both frameworks of FL and SL techniques. Section III presents the experimental results, focusing on accuracy, energy consumption, and the impact of noise and fading. Finally, Section IV provides conclusions and discusses potential directions for future work.
Figure 1: Federated learning system design
@@ II-A Federated Learning System
As shown in Fig 1, in FL, $N$ users collaborate by training local models and sharing their updates with a central server over $K$ communication cycles. The key operation in FL is the transmission of quantized model updates to minimize communication overhead and ensure data privacy, as raw data never leaves the users’ devices. As shown in Fig. 1, each communication cycle $k$ involves several key steps. Each user $i$ performs $J$ local training iterations on their private dataset to update their local model weights $W_{i}^{(k)}$. This iterative local training allows users to improve their models before sharing updates with the server. The updated local model weights $W_{i}^{(k)}$ are quantized to reduce communication overhead. The quantization process converts the weights into discrete levels, scaled by a factor $S_{i}^{(k)}$ derived from the maximum absolute weight value and the bit-width $b$ used for quantization. The quantized weights $Q_{i}^{(k)}$ are computed as:
@@
where the scale factor $S_{i}^{(k)}$ is defined as:
$S_{i}^{(k)}=\frac{\max\left(|W_{i}^{(k)}|\right)}{2^{b-1}-1}.$ (2)
After quantization, the quantized weights $Q_{i}^{(k)}$ are encoded into a bit stream $X_{i}^{(k)}$, which is then digitally modulated into a signal $Z_{i}^{(k)}$. This signal is transmitted through the wireless channel. Consequently, the received signal $\hat{Z}_{i}^{(k)}$ is a faded and noisy version of the originally transmitted signal due to Rayleigh fading $f_{i}^{(k)}$ and additive noise $n_{i}^{(k)}$. These channel effects can lead to slight changes in the demodulated bit stream $\hat{X}_{i}^{(k)}$ [13]. The server then collects the received updates $\hat{Z}_{i}^{(k)}$ from each user $i$. These signals are first demodulated into the bit stream $\hat{X}_{i}^{(k)}$, which is then decoded into the quantized weights $\hat{Q}_{i}^{(k)}$. After decoding, the server dequantizes $\hat{Q}_{i}^{(k)}$ to recover the estimated model weights $\hat{W}_{i}^{(k)}$
@@
Finally, the server aggregates these dequantized updates using Federated Averaging (FedAvg) [14] as
@@ II-B Split Learning System
Figure 2: Split learning system design for one user. The user processes the initial layers and sends intermediate activations to the server for further processing.
SL enables a distributed learning approach where the model is divided between a user and a server, as shown in Fig. 2. In each communication cycle $k$, the user processes only the initial layers of the model locally, which reduces the computational burden on the user and helps protect raw data privacy. Then, it transmits the intermediate activations to the server for further processing. Specifically, for a given input $x_{i}$ at cycle $k$, the user computes the smashed data
@@
The server computes gradients with respect to its model weights and the activations. The gradient $\nabla_{\hat{S}_{i}^{(k)}}L^{(k)}$ is clipped to manage its magnitude and then transmitted back to the user through the wireless channel. The clipped gradient is encoded into bits, modulated using BPSK, and affected by channel impairments similar to the forward transmission. The server updates its model weights using
@@ II-C Wireless Channel
In both learning systems, semantically encoded model updates or activations (denoted as $Z$) are transmitted over wireless channels subject to Rayleigh fading and additive white Gaussian noise (AWGN). The fading and noise models are respectively defined as
$f=\sqrt{X^{2}+Y^{2}},\quad X,Y\sim N(0,1),$ (11)
$n\sim N(0,\sigma^{2}).$ (12)
The fading coefficient $f$ uniformly affects all transmitted signals $Z_{i}$, and the noisy transmission is represented by
@@
4:        User $i$:
5:        $W_{i}^{(k)}=\mathbf{\hat{W}}^{(k)}$
6:        for $j=1$ to $J$ do
7:           $W_{i}^{(k)}\leftarrow W_{i}^{(k)}-\eta\nabla L_{i}(W_{i}^{(k)})$
8:        end for
9:        $Q_{i}^{(k)}=\text{Quantize}(W_{i}^{(k)},Q)$Using (1) (2)}
10:        $X_{i}^{(k)}=\text{Encode}(Q_{i}^{(k)})$
11:        $Z_{i}^{(k)}=\text{Modulate}(X_{i}^{(k)})$
12:        Transmit: $\tilde{Z}_{i}^{(k)}=f_{i}\cdot Z_{i}^{(k)}+n_{i}$, using (11) (12) (13)
13:     end for
14:     Server:
15:     $\hat{Z}_{i}^{(k)}=\text{Demodulate}(\tilde{Z}_{i}^{(k)})$ $\forall i$
16:     $\hat{X}_{i}^{(k)}=\text{Decode}(\hat{Z}_{i}^{(k)})$ $\forall i$
17:     $\hat{Q}_{i}^{(k)}=\text{Dequantize}(\hat{X}_{i}^{(k)})$ $\forall i$
18:     Obtain $\bar{W}^{(k+1)}$ using Using (4) and broadcast it to back all users.
19:  end for
@@
6:        Transmit $\tilde{Z}_{i}^{(k)}$Encoding, Modulation; Using (11), (12), (13)}
@@
14:     Transmit gradient $\tilde{Z}_{\nabla i}^{(k)}$Encoding, Modulation; Using (11), (12), (13)}
@@ II-D Communication Energy Calculation
In our approach, communication energy is determined by calculating the channel capacity $C$, which represents the maximum data rate for error-free transmission. Using the Shannon-Hartley theorem [15], the channel capacity incorporates both the Signal-to-Noise Ratio (SNR) and the bandwidth $B$. The time required to transmit one bit and the corresponding energy consumption per bit is derived from the following equations. The SNR, representing the ratio of signal power $P$ to noise power $\sigma^{2}$, is given by
$\text{SNR}=\frac{P}{\sigma^{2}}.$ (14)
Here, $P$ is the transmission power and $\sigma^{2}$ is the noise variance as defined in (12). These parameters are utilized in both FL and SL systems. Next, the channel capacity $C$ is computed as
$C=B\log_{2}(1+\text{SNR}).$ (15)
Finally, the energy consumed per transmitted bit is given by
Energy per $\text{Energy per Bit}=\frac{P}{C}\ \ J/b,$ (16)
which is multiplied by the total number of bits (payload) to estimate the consumed communication energy.
@@ III Experiments and Results
We conducted our experiments using the Sentiment140 dataset [16, 17], which contains 1.6 million tweets. To adapt the dataset for resource-constrained devices, only the text and target labels (0 for negative sentiment and 1 for positive sentiment) were retained, and the dataset size was halved. All experimental configurations are summarized in Table I. Specifically, 90% of the data was allocated for training and 10% for testing. The vocabulary was restricted to the 10,000 most frequent words, and the maximum sequence length was set to 30. We employed 3 users in the FL setup and 1 global, undergoing 7 communication cycles with a batch size of 512. user models were trained for 5 epochs per user per cycle. Binary phase shift keying (BPSK) is the digital modulation scheme adopted to transmit the quantized data. Energy consumption and CO2 emissions were monitored using the Eco2AI framework [18], with measurements taken every 10 seconds during the FL cycles.
@@ III-A Model Architecture
The model, with a total of 89,673 parameters and an approximate size of 175.14 KB using 16-bit precision, is tailored for resource-constrained environments. This compact design, combined with our dual-approach framework, offers significant benefits in resource utilization and privacy protection.
@@ III-A1 Federated learning
The FL architecture consists of several layers: an input layer that accepts fixed-length sequences of maximum length $N$, an embedding layer that transforms input sequences into dense vectors of size 8, a convolutional layer with 32 filters and a kernel size of 3. After that comes batch normalization and max pooling, an LSTM layer with 32 units to capture temporal dependencies in the sequential output, a dense layer with 16 units using ReLU activation and L2 regularization, and finally, an output layer with 1 unit using sigmoid activation for binary classification.
@@ III-A2 Split learning
In SL the model is divided between the user and the server to reduce computational demands. The user-side computation is represented by the function $f_{u}(x)$, where $x$ is the input data processed through the initial layers of the model, including convolutional and pooling layers. The output of $f_{u}(x)$ is then sent to the server, where the server-side computation $f_{s}(f_{u}(x))=y$ processes the data further using the LSTM and subsequent layers to produce the final output $y$. This split allows the server to handle the more computationally intensive tasks, thereby completing the training process efficiently.
@@ III-B Experimental Results
As shown in Fig. 3(a), both CL and FL with 8-bit (Q8) and 32-bit (Q32) quantization stabilize around an accuracy of 0.78. However, introducing noise with an SNR level of 20 dB and fading in CL as shown in Fig. 3(d) causes a slight degradation in accuracy. This degradation occurs because CL transmits raw data, which is directly impacted by noise and fading during transmission, leading to higher error rates. In contrast, FL transmits quantized model weights, which are less vulnerable to transmission impairments due to their smaller size and structured form. This enables FL to maintain higher robustness. We also assess the impact of different quantization levels (Q4, Q8, Q16, and Q32) on model performance as shown in Fig. 3(b). Lower-quantization levels (i.e., higher compression ratio) such as Q4 significantly reduce accuracy due to the loss of precision. In contrast, Q8 and higher provide a reasonable trade-off between accuracy and communication efficiency. Among the quantization levels tested, Q8 emerges as the optimal choice for FL, balancing accuracy and communication efficiency.
TABLE II: Accuracy, computation and communication energy, carbon footprints, and total bits transferred per cycle
@@
Algorithm Accuracy Comp. Energy CO2 Emissions Comm. Energy Total Bits
(J) (g) (J) (M bits)
Central 0.7803 0 21.3 2.63 115.7
FL Q8 0.7806 60.82 9.64 $\textbf{1.64}\boldsymbol{\times 10^{-3}}$ 0.72
SL (Early Cut) 0.7813 2.95 0.47 23.5 10321.9
@@
Note: Total Bits are reported per user. For all entries, computational energy is reported for the user side.
@@
In Fig. 3(c), we analyze the effect of varying SNR on model accuracy across FL, CL, and SL. Accuracy improves steadily with increasing SNR, particularly between 0 dB and 10 dB, with FL achieving the highest overall accuracy. Beyond 20 dB, accuracy plateaus at about 0.78 across all methods, with 20 dB identified as a reasonable balance between accuracy and communication power. Again, FL is more robust in noisy and fading environments. Meanwhile, Fig. 3(d) elucidates the effect of Rayleigh fading and noise together. Herein, we set the SNR to 20 dB. We observe that both FL with and SL with Q8 maintain high accuracy despite the challenging conditions, demonstrating their robustness in real-world scenarios, despite the communication channel imperfections.
Finally, Table II summarizes each approach’s user-side energy usage and CO2 emissions. Specifically, SL and FL were evaluated at a signal-to-noise ratio (SNR) of 20 dB, as indicated in the last three entries of the table. It is important to note that the table presents only the computational energy consumed on the user side. Consequently, CL appears with zero user-side computational energy consumption. However, CL incurs substantial energy usage on the server side, which is not reflected in this table.
While SL is highly efficient in terms of user-side computational power, utilizing only 3 J compared to CL’s 134 J on the server side and 61 J for FL, SL incurs the highest communication energy costs due to the frequent transmission of intermediate activations. FL achieves a balance between communication and computing efficiency, making it ideal for contexts with limited communication resources. Since computational energy on the user side is dominant over communication energy in federated and split learning approaches, these methods consume less energy on the user side in general. Specifically, SL is the most user-side energy-efficient and environment friendly in terms of CO2 emissions making it an attractive TinyML solution.
IV Endline

### v2

# TinyML NLP Scheme for Semantic Wireless Sentiment Classification with Privacy Preservation

@@ Abstract
Natural Language Processing (NLP) operations, such as semantic sentiment analysis and text synthesis, often raise privacy concerns and demand significant on-device computational resources. Centralized learning (CL) on the edge provides an energy-efficient alternative but requires collecting raw data, compromising user privacy. While federated learning (FL) enhances privacy, it imposes high computational energy demands on resource-constrained devices. This study provides insights into deploying privacy-preserving, energy-efficient NLP models on edge devices. We introduce semantic split learning (SL) as an energy-efficient, privacy-preserving tiny machine learning (TinyML) framework and compare it to FL and CL in the presence of Rayleigh fading and additive noise. Our results show that SL significantly reduces computational power and CO2 emissions while enhancing privacy, as evidenced by a fourfold increase in reconstruction error compared to FL and nearly eighteen times that of CL. In contrast, FL offers a balanced trade-off between privacy and efficiency. This study provides insights into deploying privacy-preserving, energy-efficient NLP models on edge devices.
@@ I Introduction
Artificial Intelligence (AI) has gained widespread adoption, with applications ranging from text and image classification to text and image generation. NLP has seen rapid growth, driving advancements in virtual assistants and language classification systems. Transformer-based Large Language Models (LLMs), such as OpenAI’s GPT series [1], and BERT, have significantly advanced NLP by enabling machines to perform complex tasks with high accuracy. However, these models require substantial computational resources for training and inference, making them less practical for deployment on resource-constrained devices. Additionally, older temporal models like long short-term memory networks (LSTMs) and other recurrent neural network (RNN) variants remain a popular choice due to their ability to balance computational efficiency and performance.
@@
Firstly, although more efficient than LLMs, temporal models still require substantial storage and processing power. The high dimensionality of language data and large vocabularies further amplify the computational demands [2]. Secondly, privacy remains a major concern, as training robust models often necessitates diverse datasets, potentially exposing sensitive information. CL exacerbates this risk by collecting raw data, whereas FL mitigates it by decentralizing training. However, FL remains vulnerable to inference attacks on gradient updates. SL enhances privacy by transmitting only intermediate activations, significantly reducing the risk of data reconstruction. Thirdly, deploying models on edge devices introduces challenges, including limited computational power, bandwidth constraints, and latency issues. Wireless data transmission over channels such as WiFi is particularly susceptible to noise, fading, and unstable connections, all of which can degrade communication efficiency.
@@
TinyML mixed models have evolved to address the above concerns. For instance, model compression techniques, such as pruning [3], quantization [4], and knowledge distillation [5], help reduce model sizes and computational requirements. TinyBERT [6], for example, utilizes knowledge distillation to create smaller models, while quantization techniques, like in LLaMA [7], enable efficient execution on low-resource devices. FL offers a decentralized model training approach, allowing multiple users to train a global model while keeping their data local collaboratively. To this end, SL [8] divides the model training process between users and a server. Users transmit activations from the initial model layers, reducing computational load while further limiting raw data exposure.
While TinyFedTL [9] introduced FL for resource-constrained microcontrollers and TDMiL [10] designed a framework for distributed learning in microcontroller networks, both approaches fall short of addressing key challenges in decentralized NLP. TinyFedTL focuses on federated transfer learning, omitting the exploration of alternative learning paradigms like SL and lacks considerations for wireless channel effects. Similarly, TDMiL primarily addresses computational variability and synchronization but relies on wired setups and overlooks the unique challenges of over-the-air communication. In contrast, our work bridges these gaps by proposing and evaluating decentralized learning frameworks for text emotion classification. Moreover, the extraction of information semantics and transmitting it instead of the information itself would significantly save bandwidth and reduce energy consumption [11].
Unlike TinyFedTL and TDMiL, we introduce realistic wireless channel impairments, such as Rayleigh fading and additive noise, into the learning process. This inclusion allows for a more comprehensive evaluation of decentralized learning in real-world scenarios. Furthermore, our semantic SL-based approach significantly reduces computational power and CO2 emissions by using a partitioned model architecture, with only the initial layers processed on the device. We further assess privacy preservation across FL, SL, and CL by measuring reconstruction loss, demonstrating SL’s superior resistance to data leakage. These contributions not only extend the applicability of decentralized learning to wireless NLP applications but also align with TinyML initiatives by enhancing resource efficiency and sustainability.
@@
The remainder of this paper is structured as follows: Section II covers the system design and experimental setup, including both frameworks of FL and SL techniques. Section III presents the experimental results, focusing on accuracy, energy consumption, privacy evaluation, and the impact of noise and fading. Finally, Section IV provides conclusions and discusses potential directions for future work.
Figure 1: Federated learning cycle: users train on local data, send model updates to server for aggregation
@@ II-A Federated Learning System
As shown in Fig 1, in FL, $N$ users collaborate by training local models and sharing their updates with a central server over $K$ communication cycles. The key operation in FL is the transmission of quantized model updates to minimize communication overhead and ensure data privacy, as raw data never leaves the users’ devices. In each communication cycle $k$, user $i$ performs $J$ local training iterations on their private dataset to update model weights $W_{i}^{(k)}$. This iterative local training allows users to improve their models before sharing updates with the server. The updated local model weights $W_{i}^{(k)}$ are quantized to reduce communication overhead. The quantization process converts the weights into discrete levels, scaled by a factor $S_{i}^{(k)}$ derived from the maximum absolute weight value and the bit-width $b$ used for quantization. The quantized weights are computed as
@@
where the scale factor $S_{i}^{(k)}$ is defined as $S_{i}^{(k)}=\frac{\max\left(|W_{i}^{(k)}|\right)}{2^{b-1}-1}$.
The quantized weights $Q_{i}^{(k)}$ are then encoded into a bit stream $X_{i}^{(k)}$, which is then digitally modulated into a signal $Z_{i}^{(k)}$. This signal is transmitted through the wireless channel. Consequently, the received signal $\hat{Z}_{i}^{(k)}$ is a faded and noisy version of the originally transmitted signal due to Rayleigh fading $f_{i}^{(k)}$ and additive noise $n_{i}^{(k)}$. These channel effects can lead to slight changes in the demodulated bit stream $\hat{X}_{i}^{(k)}$ [12]. The server then collects the received updates $\hat{Z}_{i}^{(k)}$ from each user $i$. These signals are first demodulated into the bit stream $\hat{X}_{i}^{(k)}$, which is then decoded into the quantized weights $\hat{Q}_{i}^{(k)}$. After decoding, the server dequantizes $\hat{Q}_{i}^{(k)}$ to recover the estimated model weights $\hat{W}_{i}^{(k)}$
@@
To obtain the global model update, the server aggregates these dequantized updates using Federated Averaging (FedAvg) [13] as
@@ II-B Split Learning System
Figure 2: Split learning system design for one user. The user processes layers and sends activations to the server for processing.
SL enables a distributed learning approach where the model is divided between a user and a server, as shown in Fig. 2. In each communication cycle $k$, the user processes only the initial layers of the model locally, which reduces the computational burden and protects raw data privacy. Then, it transmits the intermediate activations to the server for further processing. Specifically, for a given input $x_{i}$ at cycle $k$, the user computes the smashed data
@@
The server computes gradients with respect to its model weights and the activations. The gradient $\nabla_{\hat{S}_{i}^{(k)}}L^{(k)}$ is clipped to manage its magnitude and then transmitted back to the user through the wireless channel. The clipped gradient is encoded into bits, modulated, and subjected to channel impairments, similar to the forward transmission. The server updates its model weights using
@@ II-C Wireless Channel
In both learning systems, semantically encoded model updates or activations (denoted as $Z$) are transmitted over wireless channels subject to Rayleigh fading and additive white Gaussian noise (AWGN) $n$. The fading coefficient $f$ uniformly affects all transmitted signals $Z_{i}$, and the noisy transmission is represented by
@@
4:        User $i$: $W_{i}^{(k)}=\mathbf{\hat{W}}^{(k)}$
5:        for $j=1$ to $J$ do
6:           $W_{i}^{(k)}\leftarrow W_{i}^{(k)}-\eta\nabla L_{i}(W_{i}^{(k)})$
7:        end for
8:        $Q_{i}^{(k)}=\text{Quantize}(W_{i}^{(k)},Q)$Using (1)}
9:        $X_{i}^{(k)}=\text{Encode}(Q_{i}^{(k)})$
10:        $Z_{i}^{(k)}=\text{Modulate}(X_{i}^{(k)})$
11:        Transmit: $\tilde{Z}_{i}^{(k)}=f_{i}\cdot Z_{i}^{(k)}+n_{i}$, using (10)
12:     end for
13:     Server:
14:     $\hat{Z}_{i}^{(k)}=\text{Demodulate}(\tilde{Z}_{i}^{(k)})$ $\forall i$
15:     $\hat{X}_{i}^{(k)}=\text{Decode}(\hat{Z}_{i}^{(k)})$ $\forall i$
16:     $\hat{Q}_{i}^{(k)}=\text{Dequantize}(\hat{X}_{i}^{(k)})$ $\forall i$
17:     Obtain $\bar{W}^{(k+1)}$ using Using (3) and broadcast it to back all users.
18:  end for
@@
6:        Transmit $\tilde{Z}_{i}^{(k)}$Using (10)}
@@
14:     Transmit gradient $\tilde{Z}_{\nabla i}^{(k)}$Using (10)}
@@ II-D Communication Energy Calculation
In our approach, communication energy is determined by calculating the channel capacity $C$, which represents the maximum data rate for error-free transmission. Using the Shannon-Hartley theorem [14], the channel capacity incorporates the Signal-to-Noise Ratio (SNR), multiplied with the Rayleigh fading $f$ and the bandwidth $B$. The time required to transmit one bit and the corresponding energy consumption per bit is derived from the following equations. The SNR represents the ratio of transmission power $P$ to noise power $\sigma^{2}$. After calculating the SNR we proceed to calculate the channel capacity $C$ as
$C=B\log_{2}(1+|f|^{2}\text{SNR}).$ (11)
Finally, the energy consumed per transmitted bit is given by $\frac{P}{C}(J/b),$ which is multiplied by the total number of bits (payload) to estimate the consumed communication energy.
II-E Privacy Evaluation
We evaluate privacy by assessing the ease of reconstructing raw input data from transmitted information. The reconstruction error quantifies how much raw data can be recovered and is defined as:
$Error=\frac{1}{N}\sum_{i=1}^{N}(x_{i}-\hat{x}_{i})^{2},$ (12)
where $x_{i}$ is the original input and $\hat{x}_{i}$ is the reconstructed data. In FL, $\hat{x}_{i}$ corresponds to transmitted weights $W_{i}$, while in SL, it refers to intermediate activations $S_{i}$. To improve privacy, normalization of the data is applied to avoid value spikes that might result in reconstruction easier.
@@ III Experiments and Results
Transmission Power 1 mW
We conducted our experiments using the Sentiment140 dataset [15], which contains 1.6 million tweets. To adapt the dataset for resource-constrained devices, only the text and target labels (0 for negative sentiment and 1 for positive sentiment) were retained, and the dataset size was halved. All experimental configurations are summarized in Table I. Binary phase shift keying (BPSK) is the digital modulation scheme adopted to transmit the quantized data. Energy consumption and CO2 emissions were monitored using the Eco2AI framework [16], with measurements taken every 10 seconds during the FL cycles.
@@
To evaluate the privacy we used an autoencoder under the same experimental setup to measure reconstruction error. We introduced an adversary for each user and then averaged the reconstruction errors across users. For simplicity, the autoencoder was trained on the same dataset with direct access to the raw inputs and followed the same setup in Table I; in practical scenarios, where attackers lack such access, the reconstruction task would be substantially more challenging.
@@ III-A Model Architecture
The model, with a total of 89,673 parameters and an approximate size of 175.14 KB using 16-bit precision, is tailored for resource-constrained environments. This compact design, combined with our dual-approach framework, offers significant benefits in resource utilization and privacy protection. A compression encoder factoring by four is adopted to economize the use of resources.
@@ III-A1 Federated learning
The FL architecture consists of several layers: an input layer that accepts fixed-length sequences of maximum length $N$, an embedding layer that transforms input sequences into dense vectors of size 8, a convolutional layer with 32 filters and a kernel size of 3. After that comes an LSTM layer with 32 units to capture temporal dependencies in the sequential output, a dense layer with 16 units using ReLU activation and L2 regularization, and finally, an output layer with 1 unit using sigmoid activation for binary classification.
@@ III-A2 Split learning
In our SL model, the user-side computation is represented by the function $f_{u}(x)$, where $x$ is the input data processed through the initial layers of the model, including convolutional and pooling layers. The output of $f_{u}(x)$ is then sent to encoder for compression then sent to the server, where the server-side computation $f_{s}(f_{u}(x))=y$ takes input and pass it to a decoder for decompression then processes the data further using the LSTM and subsequent layers to produce the final output $y$. This split allows the server to handle the more computationally intensive tasks.
@@ III-B Experimental Results
Fig. 3(a), shows that both CL and FL with 8-bit (Q8) and 32-bit (Q32) quantization converge at approximately 0.78 accuracy. However, introducing noise with an SNR level of 20 dB and fading in CL as shown in Fig. 3(d) causes a slight degradation in accuracy. This degradation occurs because CL transmits raw data, which is directly impacted by noise and fading during transmission. In contrast, FL transmits quantized model weights, which are smaller and well-structured, making them less vulnerable to transmission impairments. As shown in Fig. 3(b), lower quantization levels (e.g., Q4) has less accuracy due to precision loss, while Q8 and higher offer a good trade-off between accuracy and communication efficiency. Q8 stands out as the optimal choice, balancing both effectively.
TABLE II: Performance comparison of decentralized learning algorithms: average metrics over 10 runs
@@
Algorithm Total Bits Accuracy Recon. Error Comp. Energy Comm. Energy Total Energy
(M bits) (J) (J) (J)
Central 115.7 0.7803 0.0154 0 0.3459 0.3459
FL Q8 0.72 0.7806 0.0671 60.8200 0.0021 60.8221
SL (Early Cut) 2580.48 0.7800 0.2681 3.4512 7.7162 11.1674
@@
Note: Total Bits are reported per user. For all entries, computational energy is reported on the user side.
@@
In Fig. 3(c), we analyze the effect of varying SNR on model accuracy across FL, CL, and the proposed SL framework. As SNR increases, accuracy improves, especially between 0 dB and 10 dB, with FL demonstrating the highest overall performance. Beyond 20 dB, accuracy plateaus at about 0.78 across all methods, with 20 dB identified as a reasonable balance between accuracy and communication power. Again, FL is more robust in noisy and fading environments. Meanwhile, Fig. 3(d) shows the effect of Rayleigh fading and noise together. We set the SNR to 20 dB and observe that both FL with Q8 and SL maintain high accuracy, demonstrating robustness under challenging real-world conditions, despite communication channel imperfections.
Table II elucidates a comprehensive comparison of the three approaches. The evaluation reveals several key insights about privacy-computation tradeoffs. The proposed SL scheme demonstrates the highest reconstruction error of 0.2681, indicating stronger privacy preservation as it becomes increasingly difficult to recover raw data from the intermediate activations $S_{i}$. In contrast, FL and CL show concerning privacy vulnerabilities with a low reconstruction error of 0.0671 and 0.0154, respectively.
Regarding user-side computational efficiency, SL and FL were evaluated at 20 dB SNR, while CL shows zero user-side computational energy in the table. Though not shown here, CL’s substantial server-side energy consumption should be considered for a complete efficiency assessment. SL requires the lowest computational power of 3.45 J compared to FL’s 60.82 J. It incurs the highest communication energy costs due to frequent transmission of intermediate activations, even with compression applied. FL, on the other hand, balances communication and computing efficiency, making it suitable for scenarios with limited communication resources. In both federated and split learning approaches, computational energy on the user side dominates over communication energy, resulting in generally lower energy consumption for the proposed TinyML SL on the user side. SL demonstrates superior user-side energy efficiency, producing 10 times lower CO2 emissions than FL and 20 times lower than CL, as measured using the Eco2AI library. These further strengthen the position of the proposed SL framework as an ideal choice for TinyML NLP classification problems.
IV Conclusion